% notes for the july 31 meeting. Purpose: Status Report. Attendees: Jielun Sun, Britt Stephens, Sean Burns, Don Lenschow, Peggy LeMone, Joe (for Steve Oncley), and Tim (for Tony Delany). Some Items: ------------ * The Hydra system is still waiting on the LiCor 7000 and the pumps (supposed to arrive last week). * A track for tram has been tentatively laid-out. The design is starting to take form and the electronics and final cost are currently being worked on. * Tether sondes will be deployed from Aug 12-17 (one balloon) and then from 19-24 (3 balloons). The 2nd deployment dates are somewhat flexible depending on status of other instruments. (as reported by Britt). * Cal gasses have been ordered for the cross-calibration. These were supposed to already be here. Britt is going to check status. * The ATD LiCor 6251 is now set to measure fluxes. * The ATD towers are being deployed over the next few days. Equipment was moved up there on Mon-Wed to avoid conflicts with Wayne(?) who is filling cylinders with "clean" air. Tower Setup is expected to be finished by next Monday (Aug 5). * The differential GPS from LTER was on loan for last week. We can possibly reserve it for sometime in the future.